OpenStudy (anonymous):

integral x^3e^(x^2)dx integration by parts?

OpenStudy (anonymous):

yes, thrice. have fun

OpenStudy (anonymous):

would u= \[e ^{x ^{2}}\]

OpenStudy (anonymous):

oh i read it wrong. i thought it was \[\int x^3e^xdx\] sorry

OpenStudy (anonymous):

so you can integrate by parts still, but i think you only need to do it once \[u=x^2,du = xdx\]

OpenStudy (amistre64):

down the e^x^2 maybe

OpenStudy (anonymous):

so you get \[\int u e^u du\] and now parts will give answer in one step, or rather you only have to use parts once

OpenStudy (anonymous):

get \[ue^u-\int e^u du=ue^u - e^u \] and you are done

OpenStudy (anonymous):

well that is wrong i forgot the \[\frac{1}{2}\] out front

OpenStudy (anonymous):

\[u=x^2,du = 2xdx,\frac{1}{2}du=xdx\] is more like it.

OpenStudy (amistre64):

\[\int x^3e^{x^2}dx=e^{x^2}*\frac{1}{4}x^4-\int 2xe^{x^2}*\frac{1}{4}x^4\] \[\int x^3e^{x^2}dx=\frac{1}{4}x^4e^{x^2}-\frac{1}{2}\int x^3e^{x^2}dx\] \[\frac{1}{2}\int x^3e^{x^2}dx+\int x^3e^{x^2}dx=\frac{1}{4}x^4e^{x^2}\] \[\frac{3}{2}\int x^3e^{x^2}dx=\frac{1}{4}x^4e^{x^2}\] \[\int x^3e^{x^2}dx=\frac{2}{3}\left(\frac{1}{4}x^4e^{x^2}\right)\] \[\int x^3e^{x^2}dx=\frac{1}{6}x^4e^{x^2}\] maybe?

OpenStudy (amistre64):

messd that one up somewhere

OpenStudy (anonymous):

are you making this one go around in a circle to get an equation? it might work!

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

x*x^4 = x^5; not x&3 ..... thats my fauxpaux

OpenStudy (anonymous):

but you are the master of the by parts method going down the diagonal as i recall, so it should be easier to use \[u=x^2,\frac{1}{2}du=xdx\] and integrate \[\frac{1}{2}\int u e^u du\] right?

OpenStudy (anonymous):

those pesky laws of exponents...

OpenStudy (amistre64):

the table works good; but i get to the circle thing here and get issues that I have to resolve thru therapy :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

alright i think i got it

OpenStudy (anonymous):

you should end up with \[\frac{1}{2}(ue^u-e^u)\] and replace u by x^2 and that will be done

OpenStudy (anonymous):

awsomeness. Thanks

OpenStudy (anonymous):

i dont understand how you went from \[\int\limits {x^3}e ^{x ^{2}}dx = 1/6 x ^{4}e ^{x ^{2}} and then 1/2(ue ^{u}-e ^{u})\]